The weights of 1,000 children, in average, is 51kg with standard deviation of 11kg. Suppose the weights are normally distributed, how many children weigh between 31kg and 69kg


1
Expert's answer
2021-02-28T07:27:17-0500

The weights of the "1000" children are normally distributed with mean "51" kg and standard deviation "11" kg.

Let "X" be a random variable denotes the weights of the children.

Then "\mu =51" and "\sigma =11" .

Let "Z= \frac{X-\mu }{\sigma}" . Then "Z=\frac{X-51}{11}".

Here we have to find "P(31<X<69)" .

Now "P(31<X<69)=P(\frac{31-51}{11}<Z<\frac{69-51}{11})"

"=P(-1.82<Z<1.64)"

"=P(0<Z<1.64)+P(0<Z<1.82)"

"=" "(0.4495+0.4656)"

"=0.9151"

Therefore number of children's weight lie between "31" kg and "69" kg are "=(1000\u00d70.9151)=915" (approximately)